OpenStudy (anonymous):

A plane flies round-trip to Philadelphia. If flies to Philadelphia at 220 miles per hour and back home with a tailwind at 280 miles per hour. If the total trip takes 6.5 hours, how many miles does the plane fly round-trip?

OpenStudy (dumbcow):

distance = rate * time you have 2 different speeds and 2 unknown times and 1 distance to philadelphia d = 220*t1 d = 280*t2 t1 + t2 = 6.5 ---> t2 = 6.5 - t1 ---> 220t1 = 280t2 --> 220t1 = 280(6.5 -t1) solve for t1 and plug it back in to find distance "d", total miles round trip is 2*d

OpenStudy (anonymous):

thnx:)

OpenStudy (dumbcow):

yw